OpenStudy (anonymous):

Solve the system by elimination 5x-21=3y 6x=4y+28

OpenStudy (anonymous):

get them in to standard form of ax+by = c first

OpenStudy (anonymous):

@helpme1.2 Can you show me how to solve this step by step its been awhile since ive done this?

OpenStudy (anonymous):

hold on, brb

OpenStudy (anonymous):

first start by putting them in to standard form 5x-3y=21 6x-4y=28

OpenStudy (anonymous):

then try to eliminate one of the variable, in this case lets eliminate y first

OpenStudy (anonymous):

5x-3y=21---> mult by 4 6x-4y=28---> mult by 3 20x-12y= 84 18x-12y = 84------then subtract the equation ------------------- 2x = 0 divide both by 2 x = 0

OpenStudy (anonymous):

to find the Y value plug in the x value in to one of the equations and solve for y

OpenStudy (anonymous):

you with me @GMAN787 ?

OpenStudy (anonymous):

Yes, sorry I was away for a few mins or so. @helpme1.2 \

OpenStudy (anonymous):

do you think you can solve for y ? you just need to plug the x value into one of the equation and solve

OpenStudy (anonymous):

@helpme1.2 Could you help me with tha too?

OpenStudy (anonymous):

ok the x value as we know is 0 so you could choose any of the given equation you like in this case lets pic 5x-3y=21 5(0)-3y = 21 ----> simplify it 0 - 3y =21 -3y = 21----> divide both sides by -3 y = -7